# Toric separable geometries and extremal Kähler metrics

### v1

@@ Abstract.
This paper introduces the framework of (local) toric separable geometries, where separable geometries come in families, each uniquely determined by an underlying factorization structure. This unifying framework captures all known explicit Calabi-extremal toric Kähler metrics, previously constructed through diverse methods, as two distinct families corresponding to the simplest factorization structures: the product Segre and the Veronese factorization structure. Crucially, the moduli of typical factorization structures has a positive dimension, revealing an immensely rich and previously untapped landscape of toric separable geometries. The scalar curvature of toric separable geometries is computed explicitly, necessary conditions for the PDE governing extremality are derived, and new extremal metrics are obtained systematically. In particular, for a $2m$-dimensional toric separable geometry, solutions of the PDE are necessarily $m$-tuples of rational functions of one variable belonging to an at most $(m+2)$-dimensional real vector space and whose denominators are determined by the factorization structure. Toric separable geometries serve as a separation of variables technique and are well-suited for the analytic study of geometric PDEs.
@@ 1.3. Products and decomposable elements
necessarily. However, it remains unclear which choices of $a_{j}^{r}$ lead to a factorization structure. With some effort, these structures and hence $a_{j}^{r}$s can be characterised as iterative products of Veronese factorization structures as described above, where each product is formed by completely decomposable, i.e., decomposable in each slot, tensors $S$ and $T$ (see 1.3.1 and Lemma 1.3.2).
@@ 2.3. Separable Kähler geometries
In the case of the Segre-Veronese factorization structure (see 1.2.1),
@@
$\delta_{a}^{b}$ is the Kronecker delta, and if $d_{i}=1$, then $\Delta_{ir}$ is defined to be 1. In the case of the Veronese factorization structure, (2.20) yields also Vandermonde identities (see 5.0.1).
@@
Let $T\in C^{\infty}(M^{0}_{\beta},TM\otimes\mathfrak{t}^{*})$ be the angular vector fields dual to $dt$, i.e. $dt(T)=Id_{\mathfrak{t}}$ and $\langle dt,T\rangle=Id_{TM^{0}_{\beta}}$. Using (2.20) and the identity $J_{x}J_{\tau}=-id$ from 2.2.1 we find
@@ 3. Extremality equation
As outlined in the introduction, a major focus of this paper is the explicit study of Calabi’s extremality problem in the toric setting. The Euler-Lagrange equation of the Calabi functional can be expressed as the scalar curvature being a Killing potential, which, in the setting of separable Kähler geometries (2.22), is Equation 2.63. Proposition 3.1.2 proves that a solution $A_{pq}$, $q=1,\ldots,d_{p}$, $p=1,\ldots,k$, of the extremality equation consists of rational functions $A_{pq}$ whose denominators are determined solely by the factorization structure and behave uniformly in grouped slots, and whose numerators’ degrees are bound explicitly. We show that when the defining tensors decompose in grouped $p$-slots, the difference $A_{pq}(x)-A_{pr}(x)$, $q\neq r$, is an explicit rational function (in some cases the zero function), which provides an exact relation between $A_{pq}$ and $A_{pr}$. Many of these findings are summarised in 3.3.8. Lastly, in 3.3.9, we introduce a practical technique for expressing scalar curvature in momentum coordinates for a large class of separable geometries.
@@
Throughout this section we work with the separable Kähler geometry (2.22) corresponding to, unless specified otherwise, a fixed Segre-Veronese factorization structure $\varphi:\mathfrak{h}\to V^{*}$ defined in 1.2.1. Its scalar curvature and extremality equation are given by 2.6.1. By 1.2.5, we have $\Gamma_{j}\in\bigotimes^{k}_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}S^{d_{i}}W_{i}^{*}$, hence $\varphi(\beta)\in\bigotimes_{i=1}^{k}S^{d_{i}}W_{i}^{*}$ for any $\beta\in\mathfrak{h}$, and the polynomials $\langle\hat{\textbf{x}}_{j},\Gamma_{j}\rangle$ and $\langle\mu,\beta\rangle$ are symmetric in grouped slots. In particular, $deg_{x_{pq}}(\langle\hat{\textbf{x}}_{j},\Gamma_{j}\rangle)=deg_{x_{pr}}(\langle\hat{\textbf{x}}_{j},\Gamma_{j}\rangle)$ for any $q,r=1,\ldots,d_{p}$ and $j=1,\ldots,k$, and similarly for $\langle\mu,\beta\rangle$. In addition, because $\Gamma_{j}$ and $\varphi(\beta)$ are symmetric in grouped slots, if it decomposes in the $(p,q)$-th slot, then it decomposes in grouped $p$-slots.
@@ 3.2. Case of decomposable defining tensors
We assume that $\Gamma_{j}$, $j\neq p$, decompose in grouped $p$-slots, i.e., $\Gamma_{j}=ins_{p}\left((e_{j}^{p})^{\otimes d_{p}}\otimes\tilde{\Gamma}_{j}\right)$ for some $\tilde{\Gamma}_{j}$ (see 1.2.5), and study $A_{pq}$, $q=1,\ldots,d_{p}$, a part of a solution of the corresponding extremality equation:
@@ 3.3. Relating $A_{pq}$ and $A_{pr}$ of a solution when $d_{p}\geq 2$
where $a+2bx+cx^{2}$ does not have a real root and $c\neq 0$, in the space of polynomials $\tilde{A}$ (see Proposition 3.1.2) whose degree is bounded by a constant depending on the following two cases: (i) $\ell_{p}=m$ and hence $\deg\tilde{A}\leq m+2$ by Proposition 3.1.2, and (ii) $\ell_{p}<m$ with $deg\tilde{A}\leq l_{p}$, where we used Corollary 3.2.3 since (3.13) are linearly independent in this case as 3.2.4 applies here: for full details see Lemma 4.1.1 and its proof. Lemma 3.2.1 shows that for the product Segre-Veronese factorization structure and $\langle\mu,\beta\rangle$ non-constant we have $\ell_{p}=d_{p}+1=3$ if $k\geq 2$, and $\ell_{p}=2$ if $k=1$. Therefore, in case (i) we have $m=3$ for $k\geq 2$, in which case we look for polynomial solutions of degree at most five, and $m=2$ for $k=1$, in which case we look for polynomial solutions of degree at most four. Case (ii) concerns polynomial solutions of degree at most three for $m\geq 4$ and $k\geq 2$.
@@
Let $p\in\{1,\ldots,k\}$ be fixed and $\varphi:\mathfrak{h}\to V^{*}$ be a Segre-Veronese factorization structure (1.12) such that $\Gamma_{j}$, $j\neq p$, decompose in grouped $p$-slots, i.e., $\Gamma_{j}=ins_{j}\left((e_{j}^{p})^{\otimes d_{p}}\otimes\tilde{\Gamma}_{j}\right)$ for some $e_{j}^{p}\in W_{p}^{*}$ (see 1.2.5), or the Veronese factorization structure. Let $(g_{\beta},J_{\beta})$ be the separable Kähler geometry (2.22) associated with this factorization structure and $\beta\in\mathfrak{h}$, $l_{p}$ be as in Lemma 3.1.1, and $L_{p}$ be as defined below (3.8). Let $A_{ir}$, $r=1,\ldots,d_{i}$, $i=1,\ldots,k$, be a solution of the corresponding extremality equation (3.7). Then, for $d_{p}\geq 1$ we have the following.
@@
Note that the assumption of the case (1) allows $\partial_{x_{pq}}\langle\mu,\beta\rangle=0$ and that Lemma 3.3.3 shows that $\varphi(\beta)$ decomposes in grouped $p$-slots. Corollary 3.2.2 provides shapes of a solution for any $d_{p}\geq 1$, and (3.39) relates them for $d_{p}\geq 2$, subsuming 3.3.1. Cases (2) and (6) are consequences of Corollary 3.2.3 and Proposition 3.3.2. To address cases (3) and (7) note that Proposition 3.1.2 provide shapes of a solution and Proposition 3.3.2 relate them. Shapes of a solution for cases (4) and (5) follow respectively from Corollary 3.2.3 and Proposition 3.1.2, and they are related via 3.3.6 and Lemma 3.3.7. ∎
@@
where $c_{j}\in\mathbb{R}$ and $\deg\tilde{\Upsilon}_{p}=\deg\Upsilon_{p}-o_{p}\leq d_{p}$, since $\deg\Upsilon_{p}\leq L_{p}=d_{p}+o_{p}$ (see Lemma 3.2.1 for explicit degrees). In particular, if $o_{p}=0$, then $\Upsilon_{p}=\tilde{\Upsilon}_{p}$, and if $O_{p}=\{j\}$, then $c_{j}=\Upsilon_{p}(-e^{p}_{j,1}/e^{p}_{j,2})$. Using formulae from 5.0.1 we arrive at
@@ 4. Complete solution set
This section recasts all known explicit extremal toric Kähler metrics as extremal separable geometries corresponding to Segre and Veronese factorization structures, and obtains new extremal geometries by determining all extremal separable geometries corresponding to the product Segre-Veronese factorization structure. In addition, we initiate the study of extremality for a more general class of factorization structures, the decomposable ones, and as a proof of concept for the general technique developed in 3.3.9, provide examples of extremal geometries for the first non-trivial member of this class. We also explain how the projective freedom in change of coordinates found in the literature corresponds to an isomorphism of factorization structures.
@@ 4.1. Preliminary lemma
as in 3.2.4. Let $S$ be the set of those $j\in\{1,\ldots,k\}$ for which the $j$-th summand of the above sum is non-constant, or equivalently, $\beta_{j}\in ins_{j}\left(S^{d_{j}}W_{j}^{*}\otimes\langle(1,0)^{\otimes(m-d_{j})}\rangle\right)$ does not lie on $\langle(1,0)^{\otimes m}\rangle$.
@@ 4.2. Extremal twisted orthotoric geometries
The extremality equation of twisted orthotoric geometry (see Subsection 2.4 and 2.6.2) reads
@@ 4.3. Extremal twisted product of Riemann surfaces
The extremality equation of the twisted product of Riemann surfaces (see Subsection 2.4 and 2.6.2) reads
@@ 4.4. Extremal separable geometries: the standard product Segre-Veronese factorization structure
where $\deg\Upsilon_{i}\leq d_{i}$. According to 3.3.9 and formulae from 5.0.1, the left hand side of the extremality equation, up to the factor $t^{2}$, reads
@@ 4.5. Decomposable Segre-Veronese factorization structure with 2 intersection points
Now we verify that these expressions satisfy the extremality equation precisely under the conditions from this proposition. We proceed case by case and use 3.3.9 to sum the sums up.
@@
    We use an analogous computation to the one from 3.3.9. To this end we work with the partial fraction decomposition of $\Upsilon_{1}(x)/(\pi_{1}+\pi_{2}x)^{3}$, the formula

### v2

@@ Abstract.
This paper introduces the framework of (local) toric separable geometries, where toric separable Kähler geometries come in families, each uniquely determined by an underlying factorization structure. This unifying framework captures all known explicit Calabi-extremal toric Kähler metrics, previously constructed through diverse methods, as two distinct families corresponding to the simplest factorization structures: the product Segre and the Veronese factorization structure. Crucially, the moduli of typical factorization structures has a positive dimension, revealing an immensely rich and previously untapped landscape of toric separable geometries. The scalar curvature of toric separable geometries is computed explicitly, necessary conditions for the PDE governing extremality are derived, and new extremal metrics are obtained systematically. In particular, for a $2m$-dimensional toric separable geometry, solutions of the PDE are necessarily $m$-tuples of rational functions of one variable belonging to an at most $(m+2)$-dimensional real vector space and whose denominators are determined by the factorization structure. Toric separable geometries serve as a separation of variables technique and are well-suited for the analytic study of geometric PDEs.
@@ 1.3. Products and decomposable elements
necessarily. However, it remains unclear which choices of $a_{j}^{r}$ lead to a factorization structure. With some effort, these structures and hence $a_{j}^{r}$s can be characterised as iterative products of Veronese factorization structures as described above, where each product is formed by completely decomposable, i.e., decomposable in each slot, tensors $S$ and $T$ (see Definition 1.3.1 and Lemma 1.3.2).
@@ 2.3. Separable Kähler geometries
In the case of the Segre-Veronese factorization structure (see Definition 1.2.1),
@@
$\delta_{a}^{b}$ is the Kronecker delta, and if $d_{i}=1$, then $\Delta_{ir}$ is defined to be 1. In the case of the Veronese factorization structure, (2.20) yields also Vandermonde identities (see Remark 5.0.1).
@@
Let $T\in C^{\infty}(M^{0}_{\beta},TM\otimes\mathfrak{t}^{*})$ be the angular vector fields dual to $dt$, i.e. $dt(T)=Id_{\mathfrak{t}}$ and $\langle dt,T\rangle=Id_{TM^{0}_{\beta}}$. Using (2.20) and the identity $J_{x}J_{\tau}=-id$ from Remark 2.2.1 we find
@@ 3. Extremality equation
As outlined in the introduction, a major focus of this paper is the explicit study of Calabi’s extremality problem in the toric setting. The Euler-Lagrange equation of the Calabi functional can be expressed as the scalar curvature being a Killing potential, which, in the setting of separable Kähler geometries (2.22), is Equation 2.63. Proposition 3.1.2 proves that a solution $A_{pq}$, $q=1,\ldots,d_{p}$, $p=1,\ldots,k$, of the extremality equation consists of rational functions $A_{pq}$ whose denominators are determined solely by the factorization structure and behave uniformly in grouped slots, and whose numerators’ degrees are bound explicitly. We show that when the defining tensors decompose in grouped $p$-slots, the difference $A_{pq}(x)-A_{pr}(x)$, $q\neq r$, is an explicit rational function (in some cases the zero function), which provides an exact relation between $A_{pq}$ and $A_{pr}$. Many of these findings are summarised in 3.3.8. Lastly, in Remark 3.3.9, we introduce a practical technique for expressing scalar curvature in momentum coordinates for a large class of separable geometries.
@@
Throughout this section we work with the separable Kähler geometry (2.22) corresponding to, unless specified otherwise, a fixed Segre-Veronese factorization structure $\varphi:\mathfrak{h}\to V^{*}$ defined in Definition 1.2.1. Its scalar curvature and extremality equation are given by 2.6.1. By Remark 1.2.5, we have $\Gamma_{j}\in\bigotimes^{k}_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}S^{d_{i}}W_{i}^{*}$, hence $\varphi(\beta)\in\bigotimes_{i=1}^{k}S^{d_{i}}W_{i}^{*}$ for any $\beta\in\mathfrak{h}$, and the polynomials $\langle\hat{\textbf{x}}_{j},\Gamma_{j}\rangle$ and $\langle\mu,\beta\rangle$ are symmetric in grouped slots. In particular, $deg_{x_{pq}}(\langle\hat{\textbf{x}}_{j},\Gamma_{j}\rangle)=deg_{x_{pr}}(\langle\hat{\textbf{x}}_{j},\Gamma_{j}\rangle)$ for any $q,r=1,\ldots,d_{p}$ and $j=1,\ldots,k$, and similarly for $\langle\mu,\beta\rangle$. In addition, because $\Gamma_{j}$ and $\varphi(\beta)$ are symmetric in grouped slots, if it decomposes in the $(p,q)$-th slot, then it decomposes in grouped $p$-slots.
@@ 3.2. Case of decomposable defining tensors
We assume that $\Gamma_{j}$, $j\neq p$, decompose in grouped $p$-slots, i.e., $\Gamma_{j}=ins_{p}\left((e_{j}^{p})^{\otimes d_{p}}\otimes\tilde{\Gamma}_{j}\right)$ for some $\tilde{\Gamma}_{j}$ (see Remark 1.2.5), and study $A_{pq}$, $q=1,\ldots,d_{p}$, a part of a solution of the corresponding extremality equation:
@@ 3.3. Relating $A_{pq}$ and $A_{pr}$ of a solution when $d_{p}\geq 2$
where $a+2bx+cx^{2}$ does not have a real root and $c\neq 0$, in the space of polynomials $\tilde{A}$ (see Proposition 3.1.2) whose degree is bounded by a constant depending on the following two cases: (i) $\ell_{p}=m$ and hence $\deg\tilde{A}\leq m+2$ by Proposition 3.1.2, and (ii) $\ell_{p}<m$ with $deg\tilde{A}\leq l_{p}$, where we used Corollary 3.2.3 since (3.13) are linearly independent in this case as Remark 3.2.4 applies here: for full details see Lemma 4.1.1 and its proof. Lemma 3.2.1 shows that for the product Segre-Veronese factorization structure and $\langle\mu,\beta\rangle$ non-constant we have $\ell_{p}=d_{p}+1=3$ if $k\geq 2$, and $\ell_{p}=2$ if $k=1$. Therefore, in case (i) we have $m=3$ for $k\geq 2$, in which case we look for polynomial solutions of degree at most five, and $m=2$ for $k=1$, in which case we look for polynomial solutions of degree at most four. Case (ii) concerns polynomial solutions of degree at most three for $m\geq 4$ and $k\geq 2$.
@@
Let $p\in\{1,\ldots,k\}$ be fixed and $\varphi:\mathfrak{h}\to V^{*}$ be a Segre-Veronese factorization structure (1.12) such that $\Gamma_{j}$, $j\neq p$, decompose in grouped $p$-slots, i.e., $\Gamma_{j}=ins_{j}\left((e_{j}^{p})^{\otimes d_{p}}\otimes\tilde{\Gamma}_{j}\right)$ for some $e_{j}^{p}\in W_{p}^{*}$ (see Remark 1.2.5), or the Veronese factorization structure. Let $(g_{\beta},J_{\beta})$ be the separable Kähler geometry (2.22) associated with this factorization structure and $\beta\in\mathfrak{h}$, $l_{p}$ be as in Lemma 3.1.1, and $L_{p}$ be as defined below (3.8). Let $A_{ir}$, $r=1,\ldots,d_{i}$, $i=1,\ldots,k$, be a solution of the corresponding extremality equation (3.7). Then, for $d_{p}\geq 1$ we have the following.
@@
Note that the assumption of the case (1) allows $\partial_{x_{pq}}\langle\mu,\beta\rangle=0$ and that Lemma 3.3.3 shows that $\varphi(\beta)$ decomposes in grouped $p$-slots. Corollary 3.2.2 provides shapes of a solution for any $d_{p}\geq 1$, and (3.39) relates them for $d_{p}\geq 2$, subsuming Remark 3.3.1. Cases (2) and (6) are consequences of Corollary 3.2.3 and Proposition 3.3.2. To address cases (3) and (7) note that Proposition 3.1.2 provide shapes of a solution and Proposition 3.3.2 relate them. Shapes of a solution for cases (4) and (5) follow respectively from Corollary 3.2.3 and Proposition 3.1.2, and they are related via Remark 3.3.6 and Lemma 3.3.7. ∎
@@
where $c_{j}\in\mathbb{R}$ and $\deg\tilde{\Upsilon}_{p}=\deg\Upsilon_{p}-o_{p}\leq d_{p}$, since $\deg\Upsilon_{p}\leq L_{p}=d_{p}+o_{p}$ (see Lemma 3.2.1 for explicit degrees). In particular, if $o_{p}=0$, then $\Upsilon_{p}=\tilde{\Upsilon}_{p}$, and if $O_{p}=\{j\}$, then $c_{j}=\Upsilon_{p}(-e^{p}_{j,1}/e^{p}_{j,2})$. Using formulae from Remark 5.0.1 we arrive at
@@ 4. Complete solution set
This section recasts all known explicit extremal toric Kähler metrics as extremal separable geometries corresponding to Segre and Veronese factorization structures, and obtains new extremal geometries by determining all extremal separable geometries corresponding to the product Segre-Veronese factorization structure. In addition, we initiate the study of extremality for a more general class of factorization structures, the decomposable ones, and as a proof of concept for the general technique developed in Remark 3.3.9, provide examples of extremal geometries for the first non-trivial member of this class. We also explain how the projective freedom in change of coordinates found in the literature corresponds to an isomorphism of factorization structures.
@@ 4.1. Preliminary lemma
as in Remark 3.2.4. Let $S$ be the set of those $j\in\{1,\ldots,k\}$ for which the $j$-th summand of the above sum is non-constant, or equivalently, $\beta_{j}\in ins_{j}\left(S^{d_{j}}W_{j}^{*}\otimes\langle(1,0)^{\otimes(m-d_{j})}\rangle\right)$ does not lie on $\langle(1,0)^{\otimes m}\rangle$.
@@ 4.2. Extremal twisted orthotoric geometries
The extremality equation of twisted orthotoric geometry (see Subsection 2.4 and Remark 2.6.2) reads
@@ 4.3. Extremal twisted product of Riemann surfaces
The extremality equation of the twisted product of Riemann surfaces (see Subsection 2.4 and Remark 2.6.2) reads
@@ 4.4. Extremal separable geometries: the standard product Segre-Veronese factorization structure
where $\deg\Upsilon_{i}\leq d_{i}$. According to Remark 3.3.9 and formulae from Remark 5.0.1, the left hand side of the extremality equation, up to the factor $t^{2}$, reads
@@ 4.5. Decomposable Segre-Veronese factorization structure with 2 intersection points
Now we verify that these expressions satisfy the extremality equation precisely under the conditions from this proposition. We proceed case by case and use Remark 3.3.9 to sum the sums up.
@@
    We use an analogous computation to the one from Remark 3.3.9. To this end we work with the partial fraction decomposition of $\Upsilon_{1}(x)/(\pi_{1}+\pi_{2}x)^{3}$, the formula